# CoRaiS: Lightweight Real-Time Scheduler for Multi-Edge Cooperative Computing

### v1

@@ V-B1 Analysis of Conventional Test Results
CoRaiS is a lightweight model that has about 4 million learnable parameters. We trained CoRaiS on four scales: (EN=5, RN=50), (EN=5, RN=100), (EN=10, RN=50) and (EN=10, RN=100). EN and RN denote the number of edges and requests respectively. The models occupy 1511M, 1961M, 1539M, and 2057M on a single GTX 2080Ti GPU during training. We test the performance of learned policies on the same scale problems, and the results are shown in Table II.
Figure 7: Testing on 1000 large-scale instances (30 edges and 400 requests) using the CoRaiS model trained on the scale with 10 edges and 100 requests. Left: More sampling improves solution quality (gap is decreasing); Right: More sampling slightly increases the computation time.
Figure 8: A case to illustrate that CoRaiS learned to optimize scheduling through load balancing without manual intervention. In this case, five edges are homogeneous, i.e. they have same computation time estimation functions. The response times of edges to backlogs are same, i.e. $b_{E}=b_{D}=b_{C}=b_{B}=b_{A}$. Since all newly arrived requests are submitted to $e_{A}$ and will be transferred from $e_{A}$, we only depict the relative position relationship between $e_{A}$ and others in the topology graph at the upper-left part, and quantifies the distance from $e_{A}$ to others in the table at the lower-left part. The middle graph presents the loads on each edge before being scheduled. The right graph shows the results after CoRaiS scheduling. The numbers in the two graphs indicate the number of requests to be executed. The upper-right table predicts the time required for each edge to respond to the scheduled requests.
Figure 9: A case to illustrate that CoRaiS is able to perceive workload of edges while scheduling. In this case, five homogeneous edges have same computation time estimation function, but the response times of edges to backlogs satisfy $b_{E}=b_{D}=b_{C}=b_{B}<b_{A}$.
Figure 10: A case to demonstrate that CoRaiS is able to recognize heterogeneity when cooperating multiple edges. In this case, the computing performance of five heterogeneous edges follows the order $E>D>C>B>A$, and the response times of the edges to their backlogs are same.
@@
(i) Gurobi can quickly obtain the optimal solutions for small-scale problems, and the average time costs for (EN=5, RN=50), (EN=5, RN=100) and (EN=10, RN=50) are less than $1s$. However, as the problem scale becomes larger, such as (EN=10, RN=100), the solving time increases significantly, even sometimes, Gurobi spends $10s$ but only get a sub-optimal solution.
(ii) The solutions generated by Gurobi(1s, 10s) are very close. In practice, Gurobi usually uses branch and bound to solve problems. Sometimes it can quickly get a good non-optimal solution, but it will keep exploring in the search space because it wants to obtain a better solution. Therefore, it is difficult to predict the computing time required by Gurobi to get a high-quality solution.
@@
(iii) Heuristic approaches (Local and Random) take near-zero time to give a solution, but the quality is far away from the optimum.
@@
(iv) The solving time taken by CoRaiS(greedy, 100, 1k) is close to $0.005s$, which is much less than Gurobi. The gap closing to 1 shows CoRaiS successfully learns an efficient policy that helps produce high-quality solutions. Therefore, CoRaiS has potential to provide real-time and high-quality decision to support efficient operations of the multi-edge cooperative computing system.
@@
(v) CoRaiS and other three learning-based models (FC1-CoRaiS, FC2-CoRaiS and FC3-CoRaiS) adopt the same decoding strategies (greedy, sampling). The comparison results show that the two alignment mechanisms of edge/request features play important roles in promoting policy learning.
@@
(vi) Because CoRaiS spends near $0.005s$ on solving the problems, we explore the performance of Gurobi(0.005s) as well. The experimental results illustrate that it is very difficult for Gurobi to solve the problem within such a short time, even to calculate an approximate solution. Then we relax the time constraints to $0.01s$ and $0.05s$. The results generated by Gurobi(0.01 s) show that double relaxation does not improve performance much. The proportion of problems that are not solved approximately is still high. Some simple problems can obtain near-optimal solutions, but the solutions are much far from the optimum. The results generated by Gurobi(0.05s) shows 10X relaxation helps a lot, since all problems get their sub-optimal solutions. However, the practical calculation time to solve the problems is usually longer than the constraint, such as $0.0886s$ for (EN=10, RN=50) and $0.1211s$ for (EN=10, RN=100), which indicates that Gurobi cannot precisely control the solving process according to time constraints, the solving time is still uncontrollable. Moreover, even if the time constraint is relaxed by 10 times, the performance still shows significant degradation when encountering large-scale problems, such as the average gap is 1.0334 for large-scale problems (EN=10, RN=100), while the average gap is 1.0009 for small-scale problems (EN=5, RN=50). Compared with that, CoRaiS stabilized the solving time consumption at $0.005s$, meanwhile, the gap increasing from 1.0148 to 1.0297 keeps relatively stable.
V-B2 Analysis of Generalization Test Results
@@
We train CoRaiS on instances under problem scale setting(EN=10, RN=100). The learned model is directly applied into larger-scale instances. The results are presented in Table III. Gurobi(10s) takes a long time $(>2s$ for average) to compute a good solution and its time cost increases as the problem scale becomes larger, from $2.8s$ for (EN=10, RN=200) to $7.2s$ for (En=50, RN=800). Compared with Gurobi(10s), CoRaiS costs a shorter time $(<0.02s)$ to obtain a high-quality near-optimal solution by sampling, and its time cost does not increase significantly, from $7ms$ for (EN=10, RN=200) to $16ms$ for (En=50, RN=800), even though the problem scale becomes 20 times larger. We force Gurobi to provide solutions within $(<0.01s)$ and $(<0.02s)$ respectively, to check whether it can get similar performance with CoRaiS. The results presented in Table III show that Gurobi can hardly solve large-scale problems within a short time. Then we relax the time limitation to 10 times, i.e. $0.2s$, the comparison results show that only a few large-scale problems can be solved.
Sampling decoding effect  We found that sampling more from the policy generated by CoRaiS can improve the solution quality while slightly increasing computation time (at $0.0001s$ level), the experimental results are presented in Fig. 7. It indicates that CoRaiS has the potential to be applied to other multi-edge cooperative computing systems, regardless of their scale, since CoRaiS is capable of generating high-quality solution in real-time through sampling.
V-B3 Analysis of Characteristic Validation Results
@@
(i) Load balancing (LB): We design five homogeneous edges and push same backlogs on them, so that the initial system-level state of edges are same, and the time of edges responding to backlogs satisfies $b_{E}=b_{D}=b_{C}=b_{B}=b_{A}$. Then we conduct 10k experiments on the system. In each experiment, 100 same requests are submitted to $e_{A}$, and CoRaiS is used to provide a scheduling solution. The results are presented in Table IV(LB). One randomly sampled case shown in Fig. 8 is used to visualize the CoRaiS scheduling results. The number of requests executed at five edges are approximately equal, and the response time are very close. Therefore, we claim that CoRaiS learned to optimize scheduling through load balancing without manual intervention.
@@
(ii) Workload perception (WP): We design five homogeneous edges and induce differences in response time to backlogs by pushing different numbers of requests to each edge, so that the initial system-level state of edges are different. The time of edges responding to backlogs satisfies $b_{E}\leq b_{D}\leq b_{C}\leq b_{B}<b_{A}$. Then 10k experiments submitting 100 requests to $e_{A}$ are conducted. CoRaiS makes decisions to schedule requests. As shown in Table IV(WP), the number of requests to be dispatched to each edge is different and follows the order $n_{E}\geq n_{D}\geq n_{C}\geq n_{B}>n_{A}$ on average. Meanwhile, there is no significant difference in response time after scheduling. Fig. 9 visualizes a case to illustrate the workload perception of CoRais. In this case, since $e_{A}$ has much heavier original workload than other edges, it keeps a small amount of the newly arrived requests and diverted most of them to speed up the completion of all requests. Therefore, we claim CoRaiS is able to perceive workload of edges while scheduling.
@@
(iii) Heterogeneity awareness (HA): We design five heterogeneous edges, that is, the computation time estimation functions of edges are different. We arrange the computing performance of edges in the order of $E>D>C>B>A$. Then we equalize the response times of the edges to backlogs by adjusting the number of requests. After that, 10k experiments are conducted. In each experiment, 100 same requests are submitted to $e_{A}$, and CoRaiS is used to provide a scheduling solution. According to the results in Table IV(HA), the more powerful edge serves more requests, and the corresponding response time of all edges are very close. We visualize a randomly sampled illustration in Fig. 10 as well. In this case, $e_{E}$ responds to the most requests while $e_{A}$ achieves the fewest, because the computation performance of $e_{E}$ is much more powerful than that of $e_{A}$. With CoRaiS scheduling, the variance of estimated responding time of such five edges is very small. Therefore, we claim CoRaiS is able to recognize heterogeneity while cooperating multi-edges.
VI Conclusion
@@
The system-level state evaluation model is introduced in this paper, to shield edges heterogeneity in hardware configuration and redefine the service capacity at edges. After that, an integer linear programming formulation is presented for the multi-edge scheduling problem. CoRaiS, a learning-based lightweight real-time scheduler, is proposed to minimize the response time over all distributed arriving requests. The experimental results demonstrate that CoRaiS successfully learned a strong policy to make high-quality scheduling in real time, irrespective of request arrival patterns and system scales.

### v2

@@ VII-B1 Analysis of Conventional Test Results
CoRaiS is a lightweight model that has about 4 million learnable parameters. We trained CoRaiS on four scales, i.e., $(N,Q,Z)\in\{(10,5,50),(10,5,100),(20,10,50),(20,10,100)\}$ where $(N,Q,Z)$ denotes the number of edges in the whole multi-edge cooperative computing system, the number of edges in an individual control region $Q$ and the number of requests $Z$. We test the performance of learned policies on the same scale problems, and the results are shown in Table III.
@@
(i) Gurobi can quickly obtain the optimal solutions for small-scale problems, and the average time costs for $(N,Q,Z)\in\{(10,5,50),(20,10,50),(20,10,100)\}$ are less than $1s$. However, as the problem scale becomes larger, such as when $(N,Q,Z)=(10,5,100)$, the solving time increases significantly, even sometimes, Gurobi spends $10s$ but only get a sub-optimal solution.
@@
(ii) Lightweight heuristic approaches (Local, Predicted and Random) take near-zero time to give a solution, but the quality is usually far away from the optimum.
@@
(iii) The solving time taken by CoRaiS(greedy, 100, 1k) is close to $0.02s$ with low variance $0.003s$, which is less than Gurobi in both Time-M and Time-S. The gap closing to 1 shows CoRaiS successfully learns an efficient policy that produces high-quality solutions. Therefore, CoRaiS has potential to provide real-time and high-quality decision to support efficient operations of the multi-edge cooperative computing system.
@@
(iv) CoRaiS and other three learning-based models (FC1-CoRaiS, FC2-CoRaiS and FC3-CoRaiS) adopt the same decoding strategies (greedy, sampling). The comparison results show that the two alignment mechanisms of edge/request features play important roles in promoting policy learning.
@@
(v) Because CoRaiS spends near $0.02s$ on solving the problems, we explore the performance of Gurobi(0.02s) as well. The experimental results illustrate that it is very difficult for Gurobi to solve the problem within such a short time, even to calculate an approximate solution. Then we relax the time constraints to $0.05s$, $1s$ and $5s$. The results generated by Gurobi(0.05 s) show that double relaxation does not improve performance much. The proportion of problems that are not solved approximately is still high. Only some simple problems can obtain near-optimal solutions. The results generated by Gurobi(1s) and Gurobi(5s) show 50X and 200X relaxation help a lot, since all problems get their sub-optimal solutions. However, the average decision-making time is too long for a real-time computing system, and the stability of time usage also performs not well.
VII-B2 Analysis of Generalization Test Results
We train CoRaiS on instances under problem scale setting $(10,5,100)$. The learned model is directly applied into medium-scale and large-scale instances. The results are presented in Table IV and Table V, respectively.
@@ VII-B2 Analysis of Generalization Test Results
In the medium-scale instances, the number of edges are limited to be more than 10 but less than 100, while the number of requests are constrained to several hundred levels. We present the results on four medium scales in Table IV, including $(20,10,200),(60,30,400),(100,50,600)$ and $(100,50,800)$. Gurobi(10s) takes a long time to compute a good solution, and its time cost increases as the problem scale becomes larger, i.e. from $2s$ for $(20,10,200)$ to $20s$ for $(100,50,800)$. However, as the scales become too large, Gurobi(10s) cannot solve the problem within the limited computing time, such as instances on $(100,50,600)$ and $(100,50,800)$. Compared with Gurobi(10s), CoRaiS costs a shorter time $(<0.05s)$ to obtain a high-quality near-optimal solution by sampling, and its time cost does not increase significantly, from $0.02s$ for $(20,10,200)$ to $0.05s$ for $(100,50,800)$, even though the problem scale becomes 20X larger. We force Gurobi to provide solutions within $0.05s$ to check whether it can get similar performance with CoRaiS. The results presented in Table IV show that Gurobi can hardly solve large-scale problems within a short time. Then we relax the time limitation to 20X and 100X, i.e. $1s$ and $5s$, the comparison results show that only a few large-scale problems can be solved. Lightweight heuristic approaches are also used as baselines. The comparison results show that even though CoRaiS has a decision-making time twice that of heuristic approaches, it can produce much better assignment decisions to realize shorter response time over all requests.
Furthermore, we conduct generalization testing on large-scale instances. The number of edges in large-scale instances is set to up to 1k, while the number of requests ranges from several thousand to 15k. Table V displays the results on five scales, including (200, 100, 2k), (600, 300, 4k), (1k, 500, 6k), (2k, 1k, 10k) and (2k, 1k, 15k). To facilitate the evaluation of different algorithms’ performance, we introduce a new metric, Cost-M, estimating the average response time over requests, measured in seconds. In the table, we observe that Gurobi(5s) and Gurobi(10s) struggle to handle the large-scale instances. Therefore, we select CoRaiS(1k) as the baseline and compare its results with lightweight heuristics. It’s evident that decision-making time increases with problem scales for all scheduling approaches. Lightweight heuristic approaches, especially Local, enable faster decision-making. However, their decision quality is significantly inferior to that of CoRaiS(1k). The Cost-M of lightweight heuristics is usually more than twice that of CoRaiS(1k). Additionally, to address the need for swift and effective decision-making in specific cases, we relax the sampling times of CoRaiS to 10, 100, and 500. The results show that with fewer samples, CoRaiS can speed up decision-making while ensuring more effective scheduling than baseline methods.
VII-B3 Analysis of Characteristic Validation Results
@@
(i) Load balancing (LB): We design five homogeneous edges and push same backlogs on them, so that the initial system-level state of edges are same, and the time of edges responding to backlogs satisfies $b_{E}=b_{D}=b_{C}=b_{B}=b_{A}$. Then we conduct 10k experiments on the system. In each experiment, 100 same requests are submitted to $e_{A}$, and CoRaiS is used to provide a scheduling solution. The results are presented in Table VI(LB). The number of requests executed at five edges are approximately equal, and the response time are very close. Therefore, we claim that CoRaiS learned to optimize scheduling through load balancing without manual intervention.
@@
(ii) Workload perception (WP): We design five homogeneous edges and induce differences in response time to backlogs by pushing different numbers of requests to each edge, so that the initial system-level state of edges are different. The time of edges responding to backlogs satisfies $b_{E}\leq b_{D}\leq b_{C}\leq b_{B}<b_{A}$. Then 10k experiments submitting 100 requests to $e_{A}$ are conducted. CoRaiS makes decisions to schedule requests. As shown in Table VI(WP), the number of requests to be dispatched to each edge is different and follows the order $n_{E}\geq n_{D}\geq n_{C}\geq n_{B}>n_{A}$ on average. Meanwhile, there is no significant difference in response time after scheduling. Therefore, we claim CoRaiS is able to perceive workload of edges while scheduling.
@@
(iii) Heterogeneity awareness (HA): We design five heterogeneous edges, that is, the computation time estimation functions of edges are different. We arrange the computing performance of edges in the order of $E>D>C>B>A$. Then we equalize the response times of the edges to backlogs by adjusting the number of requests. After that, 10k experiments are conducted. In each experiment, 100 same requests are submitted to $e_{A}$, and CoRaiS is used to provide a scheduling solution. According to the results in Table VI(HA), the more powerful edge serves more requests, and the corresponding response time of all edges are very close. Therefore, we claim CoRaiS is able to recognize heterogeneity while cooperating multi-edges.
VIII Prototype Evaluation
Figure 11: The linear relationship between running time and the size of data packets for Model Soups [37] on the five heterogeneous devices (Edge A, Edge B, Edge C, Edge D, Edge E). Points are actual statistical data, while lines are function relationships fitted based on the data.
VIII-A Environment Description
Prototype experiments are conducted to illustrate the performance of proposed CoRaiS. Through out the experiments, five heterogeneous servers are deployed as edge computing nodes, and the servers are networked using a dedicated VPN. The configurations and the computation time estimation functions (Model Soups[37] oriented) of the servers are shown in Fig. 11. The number of service replicas deployed on the edges are (2, 4, 3, 3, 1). Three levels of requests are submitted to the edges, comprising small-size requests with 2.6MB of data, medium-size requests with 7.9MB, and large-size requests with 18.6MB. Five-scale experiments are established, i.e., the number of requests involves 50, 100, 200, 500, and 1000. In each experiment, the distribution percentages for these request sizes are 50%, 30%, and 20%, respectively. The primary index collected is the physical response time, considered the most crucial metric.
Two kinds of experiments are carried out, including equal-distribution testing and shifted-distribution testing.
  • Equal-distribution testing: Each request’s source edge and predicted edge are uniformly sampled from the five available edges, while its execution edge is determined through scheduling algorithms. The distribution setting for requests aligns with the data generation method used during the training of CoRaiS.
  • Shifted-distribution testing: The probabilities of each request being submitted to the five edges are predefined, i.e. [50%, 20%, 20%, 5%, 5%], which makes the initial distribution of requests on edges be quite different from the uniform distribution. The probabilities of each request’s predicted edge is predefined as well, that is, [10%, 10%, 40%, 25%, 15%]. The execution edge of each request is decided by scheduling algorithms.
To enable real-time scheduling of requests, lightweight scheduling algorithms were utilized in the experiments. These include CoRaiS [^5] CoRaiS is trained on scheduling problems (10, 5, 100). with a sampling decoding strategy, denoted as CoRaiS(1k), as well as three heuristic approaches capable of making decisions within 0.1 seconds, i.e. Local, Predicted, and Random(1k).
VIII-B Results Analysis
For equal-distribution testing, the practical response times of requests are summarized in Table VII. From the table, it is evident that CoRaiS(1k) consistently facilitates more effective scheduling decisions, resulting in faster response times for requests.
For shifted-distribution testing, the experimental results are summarized in Table VIII. The findings outlined in Table VIII reveal that CoRaiS is adaptable to request distributions different from those it was trained on, enabling more effective decisions to accelerate response times for requests.
TABLE VII: Practical response time of requests while establishing equal-distribution testing (measured in seconds)
ReqNum CoRaiS(1k) Local Predicted Random(1k)
50 6.361 11.206 8.223 7.985
100 13.503 14.375 15.891 15.101
200 26.613 33.124 31.087 35.854
500 67.972 72.965 77.099 80.369
1000 130.385 141.639 146.179 157.577
TABLE VIII: Practical response time of requests while establishing shifted-distribution testing (measured in seconds)
ReqNum CoRaiS(1k) Local Predicted Random(1k)
50 7.454 12.834 8.624 8.187
100 13.743 22.908 17.056 16.888
200 27.604 46.715 30.434 31.749
500 67.630 115.215 94.336 78.333
1000 131.461 224.968 176.093 157.254
IX Conclusion
@@
To facilitate the effective cooperation among edges, this paper firstly introduces the system-level state evaluation model to shield edges heterogeneity in hardware configuration and redefine the service capacity at edges. Following that, the territorial schedulers are introduced into the service-oriented multi-edge cooperative computing system, and the multi-edge scheduling problem is formulated to an integer linear programming formulation to inspire the design of multi-edge cooperative algorithms. After that, CoRaiS, a learning-based lightweight real-time scheduler, is proposed to minimize the response time over all distributed arriving requests. The experimental results demonstrate that CoRaiS successfully learned a strong policy to make high-quality scheduling in real time, irrespective of request arrival patterns and system scales.
The proposed cooperation models and algorithms still have limitations in supporting multi-edge cooperation at a global level due to the lack of support for multi-scheduler cooperation. In the future, we will continue studying multi-edge cooperative computing and plan to design a hierarchical interoperability model based on current research progresses. The model will categorize schedulers into medium-level schedulers and a top-level scheduler. Medium-level schedulers can cooperate with each other under the control of their higher-level scheduler. With the improved interoperability model, scheduling problems will be reformulated from the different perspectives of medium-level schedulers and the top scheduler. Scheduling algorithms will also be designed with consideration of the accumulative decision-making time delay across the hierarchy of schedulers. The updated cooperation models and algorithms aim to ensure the effective operation of a scalable multi-edge cooperative computing system, regardless of the scheduling levels and the number of edges and requests in the system.
Acknowledgement
We are grateful to New Lossless Data Center at Purple Mountain Laboratories for generously providing sufficient computing power for model training and inferences.